# Grover Speedup from Many Forms of the Zeno Effect

### v3

@@ 1 Introduction
While definitions may vary slightly, for the purpose of the present paper by Zeno effect, we mean any interaction which prevents a quantum system from entering a state by interfering with buildup of coherences, which can be expressed mathematically as the off-diagonal elements of a density matrix. Likewise by Zeno dynamics we mean the (possibly trivial) dynamics of a system which is confined by Zeno effects.
@@
If one is to consider analog computing devices based on novel mechanisms, verifying that a mechanism has potential for a speedup is a key step to validate it as a useful tool for quantum computing. In this manuscript we focus on a speedup on unstructured search. The advantage of this setting, as opposed to general optimisation is that the scaling of the best known possible algorithm is known, so an advantage can be proven. This contrasts with combinatorial optimisation problems, where the best classical scaling is not known. Of particular relevance to this work is the recently proposed entropy computing paradigm of photonic computing [12], which aims to build analog devices making use of a number of effects, including Zeno blockade [13, 14, 15].
@@
Additionally, Zeno effects have been known to have various applications in quantum computing, for example in the aforementioned work [16]. Therein the authors demonstrate that interaction-free imaging can be extended to the few-atom level, enabling the realization of asymptotically on-demand interaction- and measurement-free quantum logic gates, which are robust against decoherence and detector inefficiency. Interaction-free Zeno gates were experimentally demonstrated in [17] building on methods proposed in [18]. Holonomic quantum gates are often proposed and implemented based on adiabatic effects to preserve a subspace [19, 20, 21]. Strong dissipation can also be used to maintain a system within a decoherence free subspace [6]. Furthermore [5] demonstrated that dissipation based Zeno gates can support universal quantum computation. More recently, [22] showed that engineered dissipative reservoirs can guide many-body quantum systems toward correlated steady states, facilitating scalable entangled many-body states and the exploration of nonequilibrium quantum phenomena. Zeno effects has also applications in error correction [23]. In [24] it was even demonstrated that a non-unitary Zeno effect implemented through measurement can be used to obtain a Grover-type advantage, a specific example of the more general result we show here. Zeno effects have also been found to be useful in implementing constraints in a gate-model setting [25].
@@
More broadly, even beyond Zeno effects, open quantum system effects can have a positive and useful effect on a system. For example improving aspects such as uniformity in quantum walks [26]. Beneficial effects from open-system interactions have also been directly used in technology, for example the reverse annealing feature as implemented in D-Wave systems only works because of finite temperature dissipation [27, 28]. For another example of the beneficial effect of open system dynamics in that setting see [29]. Consistent with these trends, our work suggests that there will be many optimal protocols which rely on open system effects but are not directly manifestations of Zeno dynamics.
@@
While real optimisation problems of interest have structure which should not be ignored, performance on unstructured search can give an indication whether it is hypothetically possible to use quantum superposition to gain an advantage. In this work we focus on whether different computational mechanisms, corresponding to different versions of Zeno effects (both continuous and discrete) can provide a quantum advantage on unstructured search. We find that for direct analog quantum computation based on all mechanisms we examine, projective measurements, phase kicks, and a form of measurement inspired by dissipation which we call destructive measurement, not only is a quantum speedup possible, but all models can reach the theoretically optimal $\sqrt{N}$ speedup, matching the scaling of the original algorithm by Grover [30, 31]. While the continuum limit of a phase kicks correspond to the adiabatic algorithm, which has already been shown to attain an optimal speedup on unstructured search [32], many of the others have not been explored. This observation highlights opportunities for novel methods of performing quantum optimisation.
@@
Unstructured search provides a natural starting place for understanding algorithms, since it operates in a relatively simple two-state subspace and has an understandable advantage. In fact unstructured search performance provided key evidence for the utility of continuous time quantum walk [33] and adiabatic quantum computing [32]. Although other evidence also played an important role, such as [34, 35] where an exponential advantage in graph traversal and some decision trees were demonstrated for continuous-time quantum walks.
@@
We structure the presentation of our work as follows. In section 2 we review the mathematical background and summarise the strategy we will use for deriving our results. Details of unstructured search Hamiltonians and the “single avoided crossing” model which we use can be found in appendices A and B. In all of these examples the systems are subject to various quantum channels. In section 2.2 we lay out the key aspects of our strategy, the conversion of the single avoided crossing model to one which is gap invariant by scaling total runtime with the gap. The next three section discuss there individual families of algorithms. In section 3 we discuss a family of algorithms based on phase rotation, where many protocols are already known, but we add some more. Section 4 discusses a family of algorithms based on decoherence-like quantum channels which is not well explored but does include search by measurement. Section 5 discusses a family of algorithms based on destruction of the excited state, something which to our knowledge has not yet been discussed in the literature. We synthesise these results into an overall picture of the three families in section 7.
@@ 2.1 Model background
The quantum channel used to solve the problem could for example be applying phases, making a projective measurement, applying decoherence, or allowing the first excited state to dissipate into a computationally useless state we call $|d\rangle$. A crucial step here is to argue over what timescales this quantum channel is physically achievable, as was done in [46] for measurements in this basis. Making arguments about the allowed timescales of these to implement the necessary quantum channels is a crucial step, if not done carefully it could lead to incorrect conclusions. For example, if we assume that dephasing or dissipation can be accomplished in constant time, independent of the size of the system, then this would lead to a system where unstructured search of any size can be accomplished in the same amount of time. Such a result can only come from an unrealistic description of a physical system: any physical quantum system can be efficiently simulated by a universal quantum computer, and it has been proven that such computers cannot exceed the $\sqrt{N}$ scaling of Grover’s algorithm [47], so such a system would have to be “unphysical” or it would contradict a known result. On the other hand, it is not a priori given that every quantum channel should result in any speedup over the classical methods. A major contribution of the current work is to show that for the models we examine, physically realistic versions are able to achieve (but not exceed) $\sqrt{N}$ scaling.
To build a Zeno effect based on discrete measurements in the present model, we must consider a continuously changing basis, which can be defined by the ground and first excited states of an unstructured search Hamiltonian. We then argue that a large but fixed (with respect to the problem size) number of repeated implementations of a quantum channel occurring at such points will maintain the original $\sqrt{N}$ scaling while faithfully describing a Zeno effect. We then further argue the existence of a continuum limit which will also retain the scaling and describe a time dependent quantum channel which implements a Zeno effect.
This work builds on a model of unstructured search known as the single avoided crossing model. The model we use was formally introduced in [36], but the underlying concept has been implicit in many other works [32, 48, 33, 49]. This model is relevant to a wide variety of cases of unstructured search, including all computationally relevant qudit models, and for both transverse field (aka hypercube) and complete graph drivers, a full discussion of the relevant search Hamiltonians appears in appendix A.
@@
Derivations of this quantity, plus a full exact solution of the single avoided crossing model appear in appendix B. A key quantity here is the energy gap between the ground and first excited state,
@@
We note that taken naively, this model appears to include terms with an infinite energy scale as $\tau\rightarrow 0$ and $\tau\rightarrow 1$. Such divergences, while they appear alarming are actually appropriate given that $f(\tau)$ is rescaled by $g_{\mathrm{min}}$. The realistic condition is that away from the avoided crossing the total energy scale $f(\tau)g_{\mathrm{min}}=O(1)$. If we applied a version which includes a cutoff
@@
the results obtained would be identical. Firstly, we can see that the proportion of $\tau$ values where this cutoff is relevant would vanish very quickly it would only be relevant when $\cot^{-1}(g^{-1}_{\mathrm{min}})/\pi>\tau$ or $\cot^{-1}(-g^{-1}_{\mathrm{min}})/\pi<\tau$. Based on the Laurante expansion of the cotangent function these correspond approximately to $\tau>g_{\mathrm{min}}$ and $\tau<1-g_{\mathrm{min}}$. Since $g_{\mathrm{min}}$ scales as $N^{-\frac{1}{2}}$, the range where this cutoff is relevant quickly vanishes. Furthermore, beyond this cutoff the Eigenstates of the system are to a very good approximation just the marked state and equal superposition. Performing a perturbative expansion, we find that in fact where $f(\tau)\approx g^{-1}_{\mathrm{min}}$, the egenstates of the system will be $\{|m\rangle+O(g^{2}_{\mathrm{min}})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(g^{2}_{\mathrm{min}})|m\rangle\}=\{|m\rangle+O(N^{-1})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(N^{-1})|m\rangle\}$. This suggests further that the error from using $f(\tau)$ instead of the more realistic $\bar{f}(\tau)$ will be of $O(N^{-1})$, which is no larger then the level of error we have already implicitly accepted from ignoring the possibility of randomly guessing the correct solution. For this reason, we can safely proceed with a model based on $f(\tau)$ despite the diverging energy values as we can be assured that the error caused by this approach could at most be as much as neglecting the possibility of random guessing.
@@ 2.2 Invariant Strategy
We base our protocols on the known optimal adiabatic schedule given in equation 3 because it provides a method to guarantee either that the protocol slows down appropriately near the crossing in the case of continuous algorithms, or that operations concentrate appropriately around the crossing in the case of protocols based on discrete operations. Note that since $H_{\mathrm{ac}}\propto g_{\mathrm{min}}$ is fixed we have implicitly set $T\propto g_{\mathrm{min}}$ (corresponding to the optimal Grover runtime), and hence do not need to directly reference it again. For discrete protocols we take evenly spaced values of $\tau$, an example of how these points are chosen is presented visually in figure 1.
Figure 1: Illustration of how an optimal annealing schedule can be used to decide parameters to perform discrete operations (measurements for example). In this example seven evenly spaced values of $\tau$ (placed at $1/8$ increments in a way which excludes 00 and $1$) are given to an optimal schedule taking the form of equation 3. These values of $f(\tau)$ then determine the values, as the top figure shows this leads to concentration around the smallest gap (emphasised with faint green line).
While all examples here arise from a quantum channel which can achieve $O(1)$ success probability on a timescale proportional to $g^{-1}(\tau)$, it is an interesting hypothetical to consider what other scaling with $g(\tau)$ could mean. Any scaling which is slower than $g^{-1}(\tau)$ would correspond to a speedup better than the known optimal (Grover) speedup, and therefore could only come about from a quantum channel which is somehow unphysical (we give an example later where an unphysical assumption about measurement accuracy appears to show such a speedup). Scaling which is faster than $g^{-1}(\tau)$, but slower than $g^{-2}(\tau)$, would correspond to a non-optimal speedup, since a quantum channel where the runtime scales as $g^{-2}(\tau)$ would result in an overall algorithmic scaling of $N$, which is the best classical scaling. Finally $g^{-2}(\tau)$ or worse would correspond to no speedup over classical scaling.
@@
As far as we are aware this is a class of algorithms which have not previously been explored in the literature. It is however reminiscent of qubit leakage models which is commonly considered in superconducting qubits [50, 51] (although relevant to any model which builds qubits from weakly anharmonic oscillators). These models however act on single physical qubits, not collective subspaces. We use the term destruction rather than dissipation to refer to this family of algorithms, to avoid confusion with computation driven by dissipation from the excited state into the ground state, an important topic, but not one which we address in the present work because it cannot be considered a source of a Zeno effect.
@@ 3.1 Applying our Strategy to Quantum Walk and Adiabatic settings
By definition the timescale of this evolution is the is the inverse gap between the ground and first excited state, which scales with $g^{-1}_{\mathrm{min}}$.
@@
It is further clear that except for very special values of $\gamma$ and $t$, the final overlap with $|m\rangle$ is non-zero, however, making sure this value does not manage as $N$ is scaled is key to demonstrate a speedup. A quantum walk at exactly the avoided crossing or within a fixed finite multiple of $g_{\mathrm{min}}$ of the avoided crossing will always attain an optimal quantum speedup, but this requires exponentially increasing precision.
@@
an optimal speedup would still occur, outside of a pathological set of cases of measure zero where the final probability of $|m\rangle$ is exactly zero. Since the subject of this paper is Zeno effects, we focus on subsets of quantum walks which have a Zeno limit. In other words, sequences where scaling up $m_{\mathrm{stage}}$ will approach a probability of $1$ of being found in the $|m\rangle$ state at the end of the protocol (as both the number of states and the total sum of scaled runtimes becomes large). Even under these restricted circumstances many such protocols exist. A natural choice is to base the spacing of the quantum walks on the optimal annealing schedule as depicted in figure 1 by setting,
@@
Figure 2: Probability of finding the ground state for a multi-stage quantum walk against $m_{\mathrm{stage}}$. Points and dotted lines are colour (or grayscale) coded based on different values of $t_{\mathrm{scale}}$). Dashed lines represent the adiabatic limit which each set of points is approaching, computed by setting $m_{\mathrm{stage}}=10,000$. The fact that the dashed lines begin at $m_{\mathrm{stage}}=50$ is an aesthetic choice and has no underlying mathematical significance.
A feature of the way in which equations 10 and 11 are defined is that the total scaled runtime $\sum^{m_{\mathrm{stage}}}_{j}t_{\mathrm{scale},j}=t_{\mathrm{scale}}$ does not depend on $m_{\mathrm{stage}}$ and therefore the runtime remains finite as $m_{\mathrm{stage}}\rightarrow\infty$. In addition to the results shown in the plots, which by themselves demonstrate an optimal advantage by virtue of being non-zero, it is straightforward to argue that an optimal advantage will be maintained in this limit. This can be done by noting that as the number of stages approaches infinity, a fixed fraction of the time spent when applying such a protocol will be between $\gamma_{\mathrm{range}}$ and $-\gamma_{\mathrm{range}}$. Specifically this fraction will be
@@
Since within this range, the eigenstates of the Hamiltonian have an overlap with both $|\omega\rangle$ and $|m\rangle$ this implies that the probability to be found in $|m\rangle$ will be non-zero starting from $|\omega\rangle$ with the possible exception of special runtimes where perfect destructive interference could occur. This proves that with the possible exceptions of special cases, these protocols will give optimal Grover speedup as well. Near the adiabatic limit the system will approximately follow the instantaneous ground state. The success probability versus $t_{\mathrm{scale}}$ in the adiabatic limit, in other words, when $m_{\mathrm{stage}}\rightarrow\infty$ is shown in figure 3.
@@
The structure of the argument makes it clear that within this rescaled picture a wide variety of schedules will yield an optimal speedup. In fact, such a schedule need only have $f(\tau)$ finite for a finite fraction of $\tau$ values, and to avoid special cases where exact destructive interference could occur. This is a more general restatement of the result of [36]. The crucial ingredient for optimal speedup is for the schedule to scale in the correct way around the avoided crossing as the problem grows other details are not particularly important.
@@
We focus on phase flipping where $\phi=\pi$. A special case is where this operation is performed on state $|\tilde{\omega}\rangle$ at the point where $f=0$. This will result in the state $|m\rangle$ with unit probability. While discrete operations of this nature seem somewhat artificial in the quantum-walk setting, the other two families we discuss will be defined starting from discrete measurement operations, so defining phase flipping, and more generally phase rotation as the primitive element of this family of algorithms is natural in the broader setting. As figure 4, shows, multiple phase rotations tend to illustrate a Zeno effect (as seen by the approach to unit marked state probability), but the dependence of the success probability on the number of phase rotation operations is far from monotonic. We also derive the Zeno effect mathematically in appendix C.

### v4

@@ 1 Introduction
Definitions of the Zeno effect and Zeno dynamics may vary slightly. For the purpose of the present paper by Zeno effect, we mean any interaction which prevents a quantum system from entering a state by interfering with buildup of coherences, which can be expressed mathematically as the off-diagonal elements of a density matrix. Likewise, by Zeno dynamics, we mean the (possibly trivial) dynamics of a system which is confined by Zeno effects.
@@
If one is to consider analog computing devices based on novel mechanisms, verifying that a mechanism has potential for a speedup is a key step to validate it as a useful tool for quantum computing. In this manuscript, we focus on a speedup on unstructured search. The advantage of this setting, as opposed to general optimisation is that the scaling of the best known possible algorithm is known, so an advantage can be proven. This contrasts with combinatorial optimisation problems, where the best classical scaling is not known. Of particular relevance to this work is the recently proposed entropy computing paradigm of photonic computing [12], which aims to build analog devices making use of a number of effects, including Zeno blockade [13, 14, 15].
Entropy computing is an optical paradigm which can support a fully quantum implementation and has demonstrated promising results in early experiments which make use of classical electronics for feedback [12]. Entropy computing is a paradigm for building photonic based optimisation machines in a way which leverages nonlinear optical effects.
Conceptually, entropy computing is similar to the coherent Ising machine paradigm [12, 16] but differs in two key aspects. Firstly, encoding is performed directly in amplitudes, as opposed to quadrature, and operation in the few-photon regime is possible. This potentially alleviates some issues with uncontrolled amplitude fluctuations disrupting an accurate quadrature encoding [16].
The second key way in which entropy computing differs from the coherent Ising paradigm is the fact that both loss and gain are controlled in the entropy computing paradigm, while only controllable gain is used in the coherent Ising setting. This controllable loss is where the Zeno blockade features in the entropy computing paradigm. By using precisely controlled fabrication techniques, non-linear optics can be used to engineer extremely high photon loss rates, which can lead to Zeno-mediated interactions even at the single photon level. Crucially, in a certain operating regime, a Zeno blockade system can exhibit the underlying effects for one of the computational models we discuss later. Because of the background required, we reserve the details for section 7.
@@
Additionally, Zeno effects have been known to have various applications in quantum computing, for example in this proposal for constructing a universal quantum computer [17]. Therein the authors demonstrate that interaction-free imaging can be extended to the few-atom level, enabling the realization of asymptotically on-demand interaction- and measurement-free quantum logic gates, which are robust against decoherence and detector inefficiency. Interaction-free Zeno gates were experimentally demonstrated in [18] building on methods proposed in [19]. Holonomic quantum gates are often designed and implemented based on adiabatic effects to preserve a subspace [20, 21, 22]. Strong dissipation can also be used to maintain a system within a decoherence free subspace [6]. Furthermore, [5] demonstrated that dissipation based Zeno gates can support universal quantum computation. More recently, [23] showed that engineered dissipative reservoirs can guide many-body quantum systems toward correlated steady states, facilitating scalable entangled many-body states and the exploration of nonequilibrium quantum phenomena. Zeno effects also have applications in error correction [24]. In [25] it was even demonstrated that a non-unitary Zeno effect implemented through measurement can be used to obtain a Grover-type advantage, a specific example of the more general result we show here. Zeno effects have also been found to be useful in implementing constraints in a gate-model setting [26].
@@
More broadly, even beyond Zeno effects, open quantum system effects can have a positive and useful effect on a system. For example, improving aspects such as uniformity in quantum walks [27]. Beneficial effects from open-system interactions have also been directly used in technology, for example the reverse annealing feature as implemented in D-Wave systems only works because of finite temperature dissipation [28, 29]. For another example of the beneficial effect of open system dynamics in that setting, see [30]. Consistent with these trends, our work suggests that there will be many optimal protocols which rely on open system effects but are not directly manifestations of Zeno dynamics.
@@
While real optimisation problems of interest have structure which should not be ignored, performance on unstructured search can give an indication whether it is hypothetically possible to use quantum superposition to gain an advantage. In this work we focus on whether different computational mechanisms, corresponding to different versions of Zeno effects (both continuous and discrete) can provide a quantum advantage on unstructured search. We find that for direct analog quantum computation based on all mechanisms we examine, projective measurements, phase kicks, and a form of measurement inspired by dissipation which we call destructive measurement a quantum speedup is possible. Furthermore, all models can reach the theoretically optimal $\sqrt{N}$ speedup, matching the scaling of the original algorithm by Grover [31, 32]. While the continuum limit of a phase kicks correspond to the adiabatic algorithm, which has already been shown to attain an optimal speedup on unstructured search [33], many of the others have not been explored. This observation highlights opportunities for novel methods of performing quantum optimisation.
@@
Unstructured search provides a natural starting place for understanding algorithms, since it operates in a relatively simple two-state subspace and has an understandable advantage. In fact, unstructured search performance provided key evidence for the utility of continuous time quantum walk [34] and adiabatic quantum computing [33]. Although other evidence also played an important role, such as [35, 36] where an exponential advantage in graph traversal and some decision trees were demonstrated for continuous-time quantum walks.
@@
We structure the presentation of our work as follows. In section 2 we review the mathematical background and summarise the strategy we will use for deriving our results. Details of unstructured search Hamiltonians and the “single avoided crossing” model which we use can be found in appendices B and C. In all of these examples, the systems are subject to various quantum channels. In section 2.2 we lay out the key aspects of our strategy, the conversion of the single avoided crossing model to one which is gap invariant by scaling total runtime with the gap. The next three section discuss their individual families of algorithms. In section 3 we discuss a family of algorithms based on phase rotation, where many protocols are already known, but we add some more. Section 4 discusses a family of algorithms based on decoherence-like quantum channels which is not well explored but does include search by measurement. Section 5 discusses a family of algorithms based on destruction of the excited state, something which to our knowledge has not yet been discussed in the literature. We synthesise these results into an overall picture of the three families in section 7.
@@ 2.1 Model background
The quantum channel used to solve the problem could for example be applying phases, making a projective measurement, applying decoherence, or allowing the first excited state to dissipate into a computationally useless state we call $|d\rangle$. A crucial step here is to argue over what timescales this quantum channel is physically achievable, as was done in [47] for measurements in this basis. Given that building on the prior work in this paper is crucial to our arguments, we review the key elements of the model from [47] in appendix A. Making arguments about the allowed timescales of these to implement the necessary quantum channels is a crucial step, if not done carefully it could lead to incorrect conclusions. For example, if we assume that dephasing or dissipation can be accomplished in constant time, independent of the size of the system, then this would lead to a system where unstructured search of any size can be accomplished in the same amount of time. Such a result can only come from an unrealistic description of a physical system: any physical quantum system can be efficiently simulated by a universal quantum computer, and it has been proven that such computers cannot exceed the $\sqrt{N}$ scaling of Grover’s algorithm [48], so such a system would have to be “unphysical” or it would contradict a known result. On the other hand, it is not a priori given that every quantum channel should result in any speedup over the classical methods. A major contribution of the current work is to show that for the models we examine, physically realistic versions are able to achieve (but not exceed) $\sqrt{N}$ scaling.
To build a Zeno effect based on discrete measurements in the present model, we must consider a continuously changing basis, which can be defined by the ground and first excited states of an unstructured search Hamiltonian. We then argue that a large but fixed (with respect to the problem size) number of repeated implementations of a quantum channel occurring within the region of parameter space where the marked state and uniform superposition both have significant overlap with the ground state will maintain the original $\sqrt{N}$ scaling while faithfully describing a Zeno effect. We then further argue the existence of a continuum limit which will also retain the scaling and describe a time dependent quantum channel which implements a Zeno effect.
This work builds on a model of unstructured search known as the single avoided crossing model. The model we use was formally introduced in [37], but the underlying concept has been implicit in many other works [33, 49, 34, 50]. This model is relevant to a wide variety of cases of unstructured search, including all computationally relevant qudit models, and for both transverse field (aka hypercube) and complete graph drivers, a full discussion of the relevant search Hamiltonians appears in appendix B.
@@
Derivations of this quantity, plus a full exact solution of the single avoided crossing model, appear in appendix C. A key quantity here is the energy gap between the ground and first excited state,
@@
We note that, taken naively, this model appears to include terms with an infinite energy scale as $\tau\rightarrow 0$ and $\tau\rightarrow 1$. Such divergences, while they appear alarming, are actually appropriate given that $f(\tau)$ is rescaled by $g_{\mathrm{min}}$. The realistic condition is that away from the avoided crossing the total energy scale $f(\tau)g_{\mathrm{min}}=O(1)$. If we applied a version which includes a cutoff
@@
the results obtained would be identical. Firstly, we can see that the proportion of $\tau$ values where this cutoff is relevant would vanish very quickly it would only be relevant when $\cot^{-1}(g^{-1}_{\mathrm{min}})/\pi>\tau$ or $\cot^{-1}(-g^{-1}_{\mathrm{min}})/\pi<\tau$. Based on the Laurante expansion of the cotangent function, these correspond approximately to $\tau>g_{\mathrm{min}}$ and $\tau<1-g_{\mathrm{min}}$. Since $g_{\mathrm{min}}$ scales as $N^{-\frac{1}{2}}$, the range where this cutoff is relevant quickly vanishes. Furthermore, beyond this cutoff the Eigenstates of the system are to a very good approximation just the marked state and equal superposition. Performing a perturbative expansion, we find that in fact where $f(\tau)\approx g^{-1}_{\mathrm{min}}$, the egenstates of the system will be $\{|m\rangle+O(g^{2}_{\mathrm{min}})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(g^{2}_{\mathrm{min}})|m\rangle\}=\{|m\rangle+O(N^{-1})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(N^{-1})|m\rangle\}$. This suggests further that the error from using $f(\tau)$ instead of the more realistic $\bar{f}(\tau)$ will be of $O(N^{-1})$, which is no larger than the level of error we have already implicitly accepted from ignoring the possibility of randomly guessing the correct solution. For this reason, we can safely proceed with a model based on $f(\tau)$ despite the diverging energy values as we can be assured that the error caused by this approach could at most be as much as neglecting the possibility of random guessing.
@@ 2.2 Invariant Strategy
We base our protocols on the known optimal adiabatic schedule given in equation 3 because it provides a method to guarantee either that the protocol slows down appropriately near the crossing in the case of continuous algorithms, or that operations concentrate appropriately around the crossing in the case of protocols based on discrete operations. Note that since $H_{\mathrm{ac}}\propto g_{\mathrm{min}}$ is fixed we have implicitly set $T\propto g_{\mathrm{min}}^{-1}$ (corresponding to the optimal Grover runtime), and hence do not need to directly reference it again. For discrete protocols we take evenly spaced values of $\tau$, an example of how these points are chosen is presented visually in figure 1.
Figure 1: Illustration of how an optimal annealing schedule can be used to decide parameters to perform discrete operations (measurements, for example). In this example, seven evenly spaced values of $\tau$ (placed at $1/8$ increments in a way which excludes 00 and $1$) are given to an optimal schedule taking the form of equation 3. These values of $f(\tau)$ then determine the values, as the top figure shows this leads to concentration around the smallest gap (emphasised with faint green line).
While all examples here arise from a quantum channel which can achieve $O(1)$ success probability on a timescale proportional to $g^{-1}(\tau)$, it is an interesting hypothetical to consider what other scaling with $g(\tau)$ could mean. Any scaling which is slower than $g^{-1}(\tau)$ would correspond to a speedup better than the known optimal (Grover) speedup, and therefore could only come about from a quantum channel which is somehow unphysical (we give an example later where an unphysical assumption about measurement accuracy appears to show such a speedup). Scaling which is faster than $g^{-1}(\tau)$, but slower than $g^{-2}(\tau)$, would correspond to a non-optimal speedup, since a quantum channel where the runtime scales as $g^{-2}(\tau)$ would result in an overall algorithmic scaling of $N$, which is the best classical scaling. Finally, $g^{-2}(\tau)$ or worse would correspond to no speedup over classical scaling.
@@
As far as we are aware, this is a class of algorithms which have not previously been explored in the literature. It is however reminiscent of qubit leakage models which are commonly considered in superconducting qubits [51, 52] (although relevant to any model which builds qubits from weakly anharmonic oscillators). These models however, act on single physical qubits, not collective subspaces. We use the term destruction rather than dissipation to refer to this family of algorithms, to avoid confusion with computation driven by dissipation from the excited state into the ground state, an important topic, but not one which we address in the present work because it cannot be considered a source of a Zeno effect.
@@ 3.1 Applying our Strategy to Quantum Walk and Adiabatic settings
By definition, the timescale of this evolution is the inverse gap between the ground and first excited state, which scales with $g^{-1}_{\mathrm{min}}$.
@@
It is further clear that except for very special values of $\gamma$ and $t$, the final overlap with $|m\rangle$ is non-zero, however, making sure this value does not vanish as $N$ is scaled is key to demonstrate a speedup. A quantum walk at exactly the avoided crossing or within a fixed finite multiple of $g_{\mathrm{min}}$ of the avoided crossing will always attain an optimal quantum speedup, but this requires exponentially increasing precision.
@@
an optimal speedup would still occur, outside of a pathological set of cases of measure zero where the final probability of $|m\rangle$ is exactly zero. Since the subject of this paper is Zeno effects, we focus on subsets of quantum walks which have a Zeno limit. In other words, sequences where scaling up $m_{\mathrm{stage}}$ will approach a probability of $1$ of being found in the $|m\rangle$ state at the end of the protocol (as both the number of states and the total sum of scaled runtimes becomes large). Even under these restricted circumstances, many such protocols exist. A natural choice is to base the spacing of the quantum walks on the optimal annealing schedule as depicted in figure 1 by setting,
@@
Figure 2: Probability of finding the ground state for a multi-stage quantum walk against $m_{\mathrm{stage}}$. Points and dotted lines are colour (or grayscale) coded based on different values of $t_{\mathrm{scale}}$. Dashed lines represent the adiabatic limit which each set of points is approaching, computed by setting $m_{\mathrm{stage}}=10,000$. The fact that the dashed lines begin at $m_{\mathrm{stage}}=50$ is an aesthetic choice and has no underlying mathematical significance.
A feature of the way in which equations 10 and 11 are defined is that the total scaled runtime $\sum^{m_{\mathrm{stage}}}_{j}t_{\mathrm{scale},j}=t_{\mathrm{scale}}$ does not depend on $m_{\mathrm{stage}}$ and therefore the runtime remains finite as $m_{\mathrm{stage}}\rightarrow\infty$. In addition to the results shown in the plots, which by themselves demonstrate an optimal advantage by virtue of being non-zero, it is straightforward to argue that an optimal advantage will be maintained in this limit. This can be done by noting that as the number of stages approaches infinity, a fixed fraction of the time spent when applying such a protocol will be between $\gamma_{\mathrm{range}}$ and $-\gamma_{\mathrm{range}}$. Specifically, this fraction will be
@@
Since within this range, the eigenstates of the Hamiltonian have an overlap with both $|\omega\rangle$ and $|m\rangle$ this implies that the probability to be found in $|m\rangle$ will be non-zero starting from $|\omega\rangle$ with the possible exception of special runtimes where perfect destructive interference could occur. This proves that with the possible exceptions of special cases, these protocols will give optimal Grover speedup as well. Near the adiabatic limit, the system will approximately follow the instantaneous ground state. The success probability versus $t_{\mathrm{scale}}$ in the adiabatic limit, in other words, when $m_{\mathrm{stage}}\rightarrow\infty$ is shown in figure 3.
@@
The structure of the argument makes it clear that within this rescaled picture, a wide variety of schedules will yield an optimal speedup. In fact, such a schedule need only have $f(\tau)$ finite for a finite fraction of $\tau$ values, and to avoid special cases where exact destructive interference could occur. This is a more general restatement of the result of [37]. The crucial ingredient for optimal speedup is for the schedule to scale in the correct way around the avoided crossing as the problem grows, other details are not particularly important.
@@
We focus on phase flipping, where $\phi=\pi$. A special case is where this operation is performed on state $|\tilde{\omega}\rangle$ at the point where $f=0$. This will result in the state $|m\rangle$ with unit probability. While discrete operations of this nature seem somewhat artificial in the quantum-walk setting, the other two families we discuss will be defined starting from discrete measurement operations. Defining phase flipping, and more generally phase rotation as the primitive element of this family of algorithms is natural in the broader setting. As figure 4, shows, multiple phase rotations tend to illustrate a Zeno effect (as seen by the approach to unit marked state probability), but the dependence of the success probability on the number of phase rotation operations is far from monotonic. We also derive the Zeno effect mathematically in appendix D.